### v1

# pnasresearcharticle

Abstract
Extreme moist heatwaves pose a serious threat to our society and human health. To manage heat-related risks, it is crucial to improve our understanding of what sets the maximum moist heat. This question has been extensively studied for moist heat in the tropics under the assumption of moist neutrality, but it remains less explored over midlatitude continents, where this assumption is not applicable. The frequent concurrent occurrence of extreme moist heat and intense convection over midlatitude land supports the prevailing hypothesis that deep convection terminates heatwaves and sets their peak intensity. Here a theory is eveloped based on the onset of convection to provide a theoretical prediction for both the maximum near-surface moist heat and potential convection over midlatitude land. Our framework demonstrates that the intensities of moist heat and convection are constrained by a preexisting energy barrier in the lower free troposphere, which suppresses convection and enables the buildup of moist heat and convective instability. This barrier, often marked by a temperature or energy inversion, varies only slightly over time, suggesting the potential to predict maximum intensities of moist heat and convection from predefined lower free tropospheric properties. This framework aligns with the stored-energy nature of deep convection over midlatitude continents, offering insights into understanding the evolution of extreme heat and convection.
keywords:
atmospheric stability moist heat deep convection convective inhibition energy inversion
Moist heatwaves, which combine the effects of high temperature and humidity, pose significant risks to human health (1). While previous research has primarily focused on the dynamics and changes of moist heat in tropical and subtropical regions (2, 3, 4), the drivers and characteristics of extreme moist heat over midlatitude continents remain poorly understood. As climate change is driving severe moist heat risks poleward into the midlatitudes (5), advancing our theoretical understanding of the limits of maximum moist heat in these regions has become increasingly urgent. Our goal in this study is to determine what sets the maximum moist heat over midlatitude land.
Near-surface wet bulb temperature (WBTs) is a key measure of moist heat stress, as it combines the influence of temperature and humidity (6, 3, 5, 7). By definition, WBT represents the temperature an air parcel would attain if cooled adiabatically to saturation at constant pressure by evaporating water into it. During this adiabatic process, the air parcel conserves its moist static energy (MSE) or the equivalent potential temperature, and thus WBTs can be directly derived assuming the conservation of near-surface MSE (MSEs). Unlike the non-linear WBT formula, MSE is a linear function of temperature, specific humidity, and geopotential height, making it simple to use. Accordingly, this study employs MSEs, defined as the sum of sensible heat, latent heat, and geopotential energy at 2 meters above the ground, to represent moist heat, to measure moist heat. The MSEs reasonably represents moist heat, as the annual maximum MSEs (Fig. 1) recovers well the spatial distribution of moist heat indicated by the annual maximum WBTs (Fig. 1). There is a small variation in WBTs given a value of MSEs, due to the dependence on the elevation variation (0β1000 m; Fig. S1). In this work, we calculated MSEs and other variables analyzed using the 3-hourly ERA5 reanalysis data during 1980β2022 for midlatitude land grids between 35βN and 75βN with elevation lower than 1000 m. Our analysis focuses on the annual maximum MSEs to explore the potential constraint on the maximum MSEs.
Regarding the maximum MSEs, previous studies have commonly assumed a strong coupling between the near-surface atmosphere and the free troposphere, such that MSEs is constrained by the free troposheric (typically 500-hPa) saturated moist static energy (MSE) through deep convection. Under the assumption of a moist-neutral atmospheric column, this framework has been widely used to study extreme heat in the tropics (3, 2). With the same neutrality assumption, recent studies (8) have applied this framework to investigate extreme dry heat over midlatitude continents, suggesting that MSEMSE. We first test this existing theory by comparing MSEs with MSE during the time of annual maximum moist heat for all land grids in the midlatitudes. Our analysis reveals that this theory does not hold for extreme moist heat, as MSEs is consistently and significantly greater than MSE in nearly all cases (Fig. 1). Furthermore, the distribution of MSE is largely zonally uniform (Fig. S1), which does not capture the strong zonal variation observed in the MSEs pattern(Fig. 1).
Here, we build on prior studies to propose a new theoretical framework that accounts for the onset of convection, which enables surface heat and convective instability to evolve beyond the neutral condition, and ultimately provides a more tight constraint on near-surface moist heat over midlatitude land.
Concurrent occurrence of moist heat and convection extremes
We first demonstrate that deep convection occurs concurrently with the annual maximum moist heat (Fig. 2), supporting the hypothesis that convective instability acts to terminate heatwaves and constrain maximum near-surface heat. The potential intensity of convection is measured by convective available potential energy (CAPE; detailed in ), with values of 1000 J kg-1 or higher typically deemed sufficient to support strong to severe convection, according to National Weather Service guidelines (9). We define the CAPE at the time of the annual maximum MSEs as critical CAPE (denoted CAPEc), which measures the potential convection that terminates the annual maximum moist heat. The CAPEc over most midlatitude land regions exceeds 1000 J kg-1, with values ranging from 3000 to 4000 J kg-1 across western Europe extending to Northeast China, and reaching 6000 J kg-1 or higher over eastern China and central North America (Fig. 2). The high values of CAPEc represent substantial convective instability in the atmosphere during extreme moist heat, indicating a significant departure from the neutrality assumption. Meanwhile, critical convective inhibition (denoted CINc), a measure of the energy barrier in the lower free troposphere that suppresses the release of CAPE (detailed in ), is sufficiently low to allow the initiation of free convection (Fig. S2).
In fact, the deep convection that terminates the maximum moist heat is likely the most intense convection of the year across many midlatitude land areas, as CAPEc (Fig. 2) closely aligns with the annual maximum CAPE (Fig. S2). This alignment is particularly evident in Europe and central North America, where the difference between CAPEc and the annual maximum CAPE is less than 250 J kg-1 (Fig. S2). In these regions, over 50% of the years between 1980 and 2022 feature simultaneous annual maxima of CAPE and MSEs (Fig. S2), emphasizing the frequent co-occurrence of extreme moist heat and intense convection in the midlatitudes.
We further provide a scaling of CAPE in terms of MSEs to demonstrate why both moist heat and convection maximize at the same time. Based on definitions of parcel buoyancy and simple assumptions about the geometry of CAPE, CAPE can be approximated by the difference between MSEs and 500-hPa saturated moist static energy (MSE) scaled by a factor of 0.22 (detailed in ):
The scaling CAPE builds a simple linear relation between CAPE and basic environmental parameters of the near-surface and free tropospheric atmospheres, without relying on the profiles of a hypothetically lifted air parcel. The scaling CAPEc, estimated using MSEs and MSE at the time of yearly maximum MSEs, effectively captures the spatial distribution of CAPEc (Fig. 2 and ), with values closely aligned along the one-to-one line (Fig. 2). We assume that MSE remains relatively constant over short time periods, which is consistent with previous studies suggesting that the free troposphere experiences minimal variation during the heatwave evolution (8) or convection build-up periods (10, 11). Consequently, based on Eq.1, CAPE are theoretically expected to evolve linearly with MSEs and reach their peaks simultaneously. The frequent concurrence of extreme near-surface moist heat and convective instability suggests a shared governing process that drives the accumulation of both phenomena over midlatitude land. This implies that the factor determining the maximum MSEs simultaneously sets the maximum CAPE. We proceed to provide a theoretical prediction for both moist heat and convection extremes.
Theoretical framework
We stick to the hypothesis that the onset of deep convection sets the maximum potential intensity of both near-surface moist heat and convection that can be achieved, but we incorporate a more detailed consideration of convection onset to propose a new theoretical framework. This framework allows surface heat and convective instability to evolve beyond the neutral condition, providing a tight constraint on both near-surface moist heat and convection throughout this process.
The foundation of our framework is the stored-energy nature of severe convection over midlatitude continents (12, 13), primarily attributed to a preexisting stable lower free troposphere (14, 10, 11). For instance, as shown in a sample sounding from a week prior to the 2014 yearly maximum surface moist heat over the central United States (temperature profiles in Fig. 3), the low-level atmospheric stability typically creates a strong layer of negative buoyancy (i.e., < 0 and CIN > 0) that inhibits convective initiation, even when the free troposphere may already be unstable (i.e., > 0 and CAPE > 0). Therefore, assuming sufficient heat sources, the near-surface atmosphere can continue to warm or moisten until the negative buoyancy layer is eroded, enabling deep free convection to initiate, release CAPE, and cool the surface. In the shown case (temperature profiles in Fig. 3), near-surface heat increased over subsequent days, accompanied by a rise in CAPE, while the free troposphere remained relatively constant, until the the time of maximum surface heat when the convective inhibition was nearly eliminated. Eventually, at the critical point when both the near-surface moist heat (MSEs) and the potential intensity of convection (CAPE) reach their maximum values, the local atmospheric environment is characterized by CINc = 0 and CAPEc 0, consistent with the results presented above (Fig. S2 and 2).
Therefore, the regime of near-surface moist heat accumulation and convection buildup is constrained by CIN 0. Since CIN essentially reflects the presence of negative buoyancy layers within the lower atmosphere, whether and when CIN approaches 0 is governed by the evolution of the most negative buoyancy () throughout the low levels. Therefore, CIN 0 is effectively equivalent to 0. The latter is more general and practical, as it also captures cases where CIN is 0 by definition due to the absence of a defined level of free convection (LFC) and equilibrium level (EL) with the entire atmospheric column characterized by negative buoyancy for the air parcel.
As parcel buoyancy is proportional to the temperature difference between the air parcel and its environment, we further utilize the conservation of the parcelβs MSE to derive an alternative formulation of temperature perturbations entirely as a function of environmental static energies (Eq.8, detailed in ). This also provides an MSE perspective to visualize the profiles of temperature perturbations and buoyancy (Eq.9, detailed in ) by comparing the parcelβs initial MSE (i.e., ) with the environmental static energy profile ( if or if , where represents the lifted condensation level). As is usually above the LCL (11), and thus based on Eq.9, 0 is equivalent to
As MSEs does not change with height, = MSEMSE, where MSE is the maximum MSEβ over the layers above or at LCL. Therefore, the maximum is limited by MSE, such that
From the equation of scaling CAPE (Eq.1), the MSE and MSE together constrain the CAPEc by
The right-hand side of Eqs.3 and 4 defines the maximum potential intensities of near-surface moist heat (MSEs) and convection (CAPE), respectively. These maximum potential intensities are achievable under conditions where surface heat sources are sufficient and external lifting forces are absent, requiring to approach zero to enable free convection, which is commonly observed in midlatitude continents (10, 11). Otherwise, the actual intensities are likely to be lower than their maximum potential values.
Maximum potential moist heat and convection
We first present the results of the sounding evolutions for the same case discussed above, analyzed within the developed MSE framework. As a validation of Eq.9, the MSE profiles (Fig. 3 and ) recover well the positive and negative buoyancy layers and their temporal evolution. Compared with temperature profiles, the MSE profiles clearly characterize the most negative buoyancy layer prior to the yearly maximum moist heat (MSE profiles in Fig. 3), identified by an explicit MSEβ inversion above the LCL. Compared to temperature profiles, the MSE profiles more clearly delineate the most negative buoyancy layer prior to the annual maximum moist heat, identified by a distinct MSEβ inversion above the LCL (Fig. 3). The preexisting MSE resides at the top of the MSEβ inversion near 800 hPa and eventually aligns with the LCL (MSE profiles in Fig. 3 and ). In this case, the final MSE effectively sets the maximum MSEs achieved at the time of the annual maximum moist heat, when (and hence CINc) approaches 0 and CAPEc is maximized. Relative to the significant changes in MSEs during this period, the preexisting MSE intensities only slightly over time, which suggests the possibility of predicting the maximum intensity of MSEs (and consequently CAPEc) based on the initial MSE, although adjustments are needed over time as MSE evolves.
To evaluate the generality of the results from the case study, we extend the analysis by presenting composite time series of key proxies derived from a broader set of cases over midlatitude land. Specifically, we focus on instances of concurrent annual maxima in MSEs and CAPE for each year during 1980β2022. To best align with our theoretical framework - where MSEs and CAPE are maximized as preexisting high CIN is eroded to zero - we further filter cases based on the following criteria: CINc = 0, CAPEc > 500 J kg-1, and a maximum CIN of at least 150 J kg-1 within the seven days preceding the critical time. These selection criteria yield 535 cases, for which we plot composite 14-day time series centered on the time of maximum MSEs (Fig. 3β). Following the maximum potential for moist heat defined by Eq.3, MSEs increases over time and ultimately converges toward MSE at the peak time (Fig. 3). Notably, MSE fails to predict the maximum MSEs as discussed above, as MSEs exceeds MSE as early as seven days before the peak, with the difference rapidly amplifying over time (Fig. 3). This growing discrepancy between MSEs and MSE reflects the progressive accumulation of CAPE leading up to the peak, as recovered by the scaling CAPE following Eq.1 and constrained by the maximum potential for convection defined in Eq.4 (Fig. 3). Fundamentally, the composite temporal evolutions of MSE and CAPE are linked to the removal of CIN or (Fig. 3), supporting our basic hypothesis formulated in Eq.2. Similarly, the composite results reveal relatively steady time series for MSE and MSE (Fig. 3), as well as for the potential CAPE (Fig. 3), highlighting the potential to predict the maximum intensities of MSEs and CAPE based on preexisting free tropospheric properties.
Lastly, to comprehensively validate the potential intensity framework (i.e., Eqs.3 and 4), we compare the observed annual maximum moist heat (MSEs) and the associated critical convection (CAPEc) with their potential maxima derived for all midlatitude land grids during 1980β2022 using the ERA5 reanalysis data (Fig. 4). Note that our goal here is not to evaluate the predictive capability of our theory based on prior atmospheric properties but rather to provide a precise, apple-to-apple assessment at the peak time. Accordingly, MSE and MSE are calculated at the time of the annual maximum MSEs. Overall, the potential MSEs (Fig. 4) and the potential CAPEc (Fig. 4) well capture the spatial patterns of the observed maximum MSEs (Fig. 1) and CAPEc (Fig. 2) over midlatitude land, with observed maxima only slightly smaller than their potential counterparts. The pattern correlation coefficient is 0.94 between observed and potential MSEs, and 0.81 between observed and potential CAPEc. The joint histograms of all cases further confirm that the potential intensities serve as a tight upper limit for the observed maximum MSEs and CAPEc (Fig. 4 and ). For the maximum MSEs, the potential MSEs is achievable for many cases, with others closely aligned with the one-to-one line but slightly below (Fig. 4). Similarly, the potential CAPEc acts as a tight constraint for the observed CAPEc as well (Fig. 4), though the variation in observed CAPEc for a given potential CAPEc appears larger, partly due to the variability introduced by the scaling of CAPEc (Fig. 2)
Conclusions and discussions
This study proposed a theoretical prediction for maximum moist heat (MSEs) and convection (CAPE) over midlatitude land, both tightly constrained by a preexisting energy (MSEβ) inversion within the lower free troposphere. The maximum MSEβ at the energy inversion layer (i.e., MSE) acts as a substantial energy barrier that suppresses the free convection of near-surface air parcels, thereby allowing the warming or moistening of near-surface air and the accumulation of convective instability. This process continues until MSEs reaches MSE, such that the near-surface air parcels are sufficiently energetic to overcome the energy barrier and initiate deep free convection. Recent studies have explored a similar topic but for dry heat extremes under the assumption of moist neutrality (8), suggesting that the 500-hPa MSE serves as an upper bound for MSEs. However, this assumption fails for extreme moist heat as MSEs generally exceeds MSE during the maximum moist heat period in midlatitudes. Notably, our scaling CAPE framework demonstrates that the surplus of MSEs relative to MSE effectively defines the potential convection intensity of the environment. This study extended prior work by carefully addressing the onset of convection, which enables the atmosphere to evolve beyond the neutral condition, consistent with the stored-energy nature of severe convection over midlatitude continents. Future work may build upon our work to incorporate other processes (e.g., wind shear, entrainments, and external forcings) to more precisely determine the onset of convection and, consequently, the termination of heatwaves. For example, recent research has highlighted the critical role of dry air entrainment in suppressing deep convection and intensifying moist heat in tropical and subtropical regions (4), though the entrainment influence might be less important in midlatitude continental convection due to the generally strong vertical wind shear (15).
The preexisting MSEβ inversion within the lower free troposphere is typically marked by a pronounced temperature inversion or a layer where the temperature decreases most gradually with height. While this study proposed the theoretical framework based on evolution of local convective instability, the formation of preexisting inversion layers might be non-local due to advection of a warm layer from the upstream region or the adiabatically warmed air from the subsidence of large-scale circulation (e.g., anticyclones). Thus, it might be necessary to investigate the origins and trajectories of the inversion layers from the Lagrangian perspective (10, 11, 16) to provide insight into the formation or potential magnitude of MSE, and thus improve the predictability of local potential moist heat and convection. This is also important in understanding the role of preexisting inversion layers in future climate, as the shift in land types or non-uniform warming by regions may affect the formation and transportation trajectories of inversion layers from region to region (17, 11).
[…]
Buoyancy and temperature perturbation
To link surface moist heat with convection, we begin by deriving an expression for parcel buoyancy as a function of near-surface moist static energy.
Considering an undiluted air parcel adiabatically lifted upward from the near-surface (2 m above ground surface; denoted by subscript ββ in derivations) and neglecting the relatively small contribution to density changes due to pressure perturbation and virtual temperature correction(18, 19), the parcel buoyancy at a given height is proportional to differences in sensible heat between the air parcel and the environment (), where is temperature perturbations of the air parcel (, with subscript ββ referring to the air parcel) with respect to the environment () and J kg-1 K-1 is the the specific heat capacity of air at constant pressure. During the adiabatic process, it is common to assume that the air parcel conserves its moist static energy (MSE) such that . is defined as the sum of sensible heat (), latent heat (), and geopotential energy (), where J kg-1 is the latent heat of vaporization, m s-2 is gravitational acceleration, is temperature, is specific humidity, and is height above sea level. Therefore, the vertical profile of is given by
where is dry static energy of the environment.
To derive solely as a function of environmental parameters, independent of the lifted air parcelβs profiles, the specific humidity of the air parcel () in Eq.5 can be further approximated using environmental parameters that vary with height () relative to the lifted condensation level (LCL).
For , the air parcel remains unsaturated and maintains its specific humidity at the initial value. Hence,
For , the air parcel is saturated, such that the parcel specific humidity equals its saturated specific humidity (, with superscript ββ referring to a quantity at saturation), which is a function of the parcel temperature and air pressure. As pressure difference between the air parcel and environment is negligible (18, 19), the difference in saturated specific humidity between the air parcel and environment is caused by difference in temperature, and thus can be written as a linear relation to the environmental saturated specific humidity () via . We further assume constant and approximate it by using Clausius-Clapeyron equation, where J kg-1 K-1 is gas constant for water vapor. Hence,
Substituting Eqs.6 and 7 into Eq.5 and rearranging it gives
Eq.8 well recovers the temperature perturbation profiles ( , Fig. S3). The denominator is always positive, ranging from 3 in the lower free troposphere above LCL to 1 in the upper free troposphere ( , Fig. S3). To simplify and derive a more intuitive form, we approximate the denominator as 1, which does not alter the sign, vertical structure, or relative magnitude of above LCL but only slightly amplifies temperature difference near LCL ( , Fig. S3). Therefore,
Eq.9 provides an alternative approach to quantify temperature perturbation profile from the MSE prospective (a more precise form is given by Eq.8; , Fig. S3). This eliminates the need to first calculate the temperature profile of a hypothetically lifted air parcel. Instead, the vertical profile of parcel buoyancy can be determined directly by comparing the parcelβs initial MSE (i.e., ) with the environmental static energy profile ( if or if ). Furthermore, this approach is equivalent to comparing near-surface potential temperature with the potential temperature profile for or the near-surface equivalent potential temperature with the saturation equivalent potential temperature profile for .
CAPE and scaling CAPE
We calculate convective available potential energy (CAPE) for the near-surface air parcel by integrating virtual temperature difference between the air parcel () and environment () with respect to natural logarithm of pressure (ln) from the level of free convection (LFC) to equilibrium level (EL), given by
where , J kg-1 K-1 is the ideal gas constant of dry air, and are pressure at LFC and EL, respectively. Moist adiabats for the air parcel follow irreversible pseudoadiabatic process (20). This makes little differences as compared to reversible process (21). Based on ideal gas law, Eq.10 is equivalent to the integral of parcel buoyancy with respect to height. The convective inhibition (CIN) is calculated in the same way as CAPE but for negative buoyancy from the surface to LFC. In this work, CIN is defined as positive by its absolute value.
Next, we derive a scaling for CAPE from Eq.10 combining simple assumptions. Since is zero at both LFC and EL and reaches its maximum approximately midway between these levels ( 500 hPa, typically above LCL), CAPE can be geometrically approximated as the area of a triangle with being the base and at 500 hPa (, with superscript ββ referring to a quantity at 500 hPa) being the height. Further neglecting virtual temperature correction (i.e., ) and substituting Eq.8, we have that
Here we focus on midlatitude extreme moist heat conditions that CAPE is in general high (1000 J kg-1). For those cases, the LFC is near the lower free troposphere and EL is close to the tropopause, and we approximate them by hPa and hPa. The mean K gives kg kg-1. Substituting these typical values into Eq.11, we provide a scaling for midlatitude high CAPE:
Previous studies have explored the relationship between CAPE and environmental MSE, either under the radiativeβconvective equilibrium framework (22, 23, 12) or through linear regression methods (24). In contrast, our derivation begins with the fundamental definitions of parcel buoyancy and CAPE, similar to (25) but incorporating minimal assumptions about the geometric characteristics of CAPE, to establish a robust linear relationship between CAPE and key atmospheric parameters near the surface and in the free troposphere. The coefficient in Eq.12 may vary with , , and , which can be precisely adjusted using the full scaling equation (Eq.11) when applied to different regions or datasets. Here, the derived coefficient of 0.22 for high-CAPE environments over midlatitude land aligns closely with the linear regression results (0.22β0.29) reported for high CAPE over continental North America (24).
This research is part of the MIT Climate Grand Challenge on Weather and Climate Extremes. Support was provided by Schmidt Sciences, LLC. We thank Kerry Emanuel, Paul OβGorman, P. J. Tuckman, and Divya Rea for helpful discussions.
[…]